# Identifying bimodality in data using a proximity-based null model with an application to classifying cell cycle phases using oxidative stress

### v2

@@ Abstract
Detecting communities in large complex networks has found a wide range of applications in physical, biological, and social sciences by identifying mesoscopic groups based on the links between individual units. Moreover, community detection approaches have been generalized to various data analysis tasks by constructing networks whose links depend on individual units’ measurements. However, identifying well-separated subpopulations in data sets, e.g., multimodality, still presents challenges as both community detection with existing null models and other partition methods give partitions that do not necessarily correspond to multimodality. Here we introduce a new spatially informed null model for this task that takes into account spatial structure but does not explicitly depend on distances between measurements. We find that community detection using this null model successfully identifies subpopulations in multimodal data and accurately does not for unimodal data. We apply this new null model to first distinguish interphase and mitotic cell cycle phases and then S and G2 cell cycle phases in a group of Dictyostelium discoideum cells using measurements of oxidative stress, which have been shown to correlate strongly with cell-cycle behaviors.
Identifying subpopulations corresponding to different behavioral or phenotypical patterns of ensembles of units using measurements of those units is a problem with a wide range of applications . One such example is identifying cell cycle phases of individual cells using oxidative stress measurements, which have been shown to correlate interphase vs mitotic phases . Using community detection in inferred networks to identify the existence or absence of multiple subpopulations presents some challenges, however, due to the generic emergence of communities in networks that are effectively spatial. In this work we present a new spatially informed null model to use with community detection via modularity to identify subpopulations in data. Using this null model, we can accurately distinguish between unimodal and multimodal data. We test this new null model over a range of synthetic benchmarks that vary separation and asymmetry in data before applying it to real data of oxidative stress in Dictyostelium discoideum cells.
I Introduction
@@
Community detection in complex networks has found applications in a wide range of applications by identifying functional and mesoscopic groups based on patterns of interactions and relationships between individual units Girvan2002PNAS ; Danon2005JStat ; Newman2006PNAS ; Fortunato2010PhysRep . Examples include groupings in online social networks Red2011SIAMRev , partnerships in political institutions Porter2005PNAS ; Porter2007PhysA ; Zhang2008PhysA ; Macon2012PhysA , architectures of neuronal organization Bassett2011PNAS ; Wymbs2012Neuron ; Yin2020PNAS , identifying antibiotic resistance in patients Parker2015ACS , and analysis of contact networks for sexually-transmitted diseases Billock2020Sex . In principle, these methods for identifying groups can be generalized to quantitative data sets by constructing networks from the data where nodes correspond to measurements and links are placed based on the pair-wise distances between measurements. These generalized methods for network-based data analysis, however, have their drawbacks, as community structures generically emerge in spatially embedded networks even when the data shows no signs of having significant subpopulations, i.e., multimodality Gilarranz2020SciRep . Thus, developing community detection frameworks that correspond more precisely with the identification of subpopulations, i.e., multimodality, and likewise accurately failing to do so in the absence of subpopulations, is an important task, as testing for multimodality in real datasets remains a challenging problem with few robust methodologies Hartigan1985AnnStats .
@@
One such problems is the classification of cell cycle phases using measurements of oxidative stress. Reactive oxygen species are hypothesized to play important roles in microbial development of cells Aguirre2005 ; Liu2020 . Reactive oxygen species are also linked with cell cycles, with mitotic cells having been shown to exhibit higher levels of oxidative stress than cells in interphase. Thus, subpopulations of cells distinguished by oxidative stress may correspond to different cell cycles. Given the practical challenges of taking individual cell measurements, linking cell cycle phases with measurements of oxidative stress could be useful in synthetic engineering tasks such as synchronizing cell cycles as well as developing a better understanding of the sources of heterogeneity between individual cells and throughout entire groups Allcroft2023 . Here we seek to identify significant subpopulations in a group of Dictyostelium discoideum cells using oxidative stress measurements quantified by the log-ratio of dihydrodichlorofluorescein diacetate (DCF) to carboxyfluorescein diacetate (CF). Technical details for the experimental measurments and data collection are provided in the Appendix.
@@
To identify subpopulations in this and other datasets we utilize the modularity method for detecting communities in networks Newman2004PRE . This method is based on comparing an actual network structure to an appropriate null model that quantifies the expectation of the existence and strength of each pairwise link. Rather than using the classical Newman-Girvan null model Newman2004PRE or other null models that incorporates spatial structure by explicitly using distances Expert2011PNAS ; Simini2012Nature ; Sarzynska2016CompNet ; Cerina2012PLOSOne , we introduce a new null model that is spatially informed but does not explicitly use distances. Thus, similar to the Newman-Girvan null model, this new null model uses primarily the local connectivity properties of nodes, but like distance-dependent null models a strong spatial structure is assumed without strong mixing. In essence, the null model is designed to assume that a node is most likely connected to the number of nodes matching its degree that are all close by, but this likelihood is not explicitly a function of any distance. We test this new null model on a series of synthetic data benchmarks to examine the effects of separation and asymmetry in the data, finding that it successfully identifies subpopulations, i.e., multimodality, when they are present and accurately fails to do so when they are not, i.e., the data is unimodal. We then turn to the task of partitioning cells according to their cell cycle phases using measurements of oxidative stress.
@@
The remainder of this paper is organized as follows. In Sec. II we describe network-based approaches to identifying subpopulations in data using modularity and discuss the choice of null models. We then introduce a new spatially informed null model designed for identifying multimodality in data. In Sec. IV we evaluate this method with our new null model over a collection of synthetic data benchmarks. In Sec. V we apply our method to real data taken from an ensemble of Dictyostelium discoideum cells. Specifically, we use oxidative stress measurements first to identify cells in interphase vs mitosis, then in S and G2 phases. In Sec. VI we conclude with a discussion of our results.
@@
II Data analysis, community detection, and the failure of existing null models
@@
Motivated by the task of identifying cell-cycle phases in experimental data of oxidative stress in Dictyostelium discoideum cells, we consider data sets $\{x_{i}\}_{i=1}^{N}$ consisting of $N$ measurements with $i=1,\dots,N$. In our motivating example, each $x_{i}$ represents a relative measurement of oxidative stress in a cell, which is strongly correlated with the phase (interphase or mitosis) of the cell cycle. (More details on the data are given in Sec. V and in Appendix A) In general we assume that each data point is a scalar, i.e., $x_{i}\in\mathbb{R}$, and therefore can be sorted in ascending order so that $x_{1}\leq x_{2}\leq\cdots\leq x_{N}$. We then construct from this data set a network of $N$ nodes where each node $i$ corresponds to a measurement $x_{i}$ and links are placed between nodes based on the distance between corresponding measurements. Many choices can be reasonably made in terms of how links are placed, but here we consider the simple choice where a link is placed between nodes $i$ and $j$ if the measurements $x_{i}$ and $x_{j}$ satisfy $|x_{i}-x_{j}|\leq rs_{x}$, where $r$ is a tunable local connectivity parameter and $s_{x}$ is the standard deviation of the dataset $\{x_{i}\}_{i=1}^{N}$. This network is described by the adjacency matrix $A$, whose entries $A_{ij}=1$ if a link exists between nodes $i$ and $j$, and otherwise $A_{ij}=0$.
@@
We then seek to map the task of identifying subpopulations in the data, i.e., multimodality, to finding communities in the corresponding network. Community detection has long been a topic of interest in the complex networks community and has a rich literature Girvan2002PNAS ; Danon2005JStat ; Fortunato2010PhysRep . Here we choose the modularity method Newman2006PNAS , where a given network structure is compared to an appropriately chosen null model and then partitioned into clusters that are more highly connected than expected by maximizing the so-called modularity. Mathematically, the modularity $Q$ can be expressed as
@@
Instead, our challenge here is in finding a null model (encapsulated in the matrix $P$) that allows us to find communities that accurately map to subpopulations, i.e., multimodality, in the original data. The classical choice for a null model was first presented by Newman and Girvan Newman2004PRE and is given by the expectation of a link being proportional to the product of the node-wise degrees, i.e., $P_{ij}=k_{i}k_{j}/(2M)$, where $k_{i}=\sum_{j=1}^{N}A_{ij}$ is the degree of node $i$. This null model is very effective when the network structure is well mixed and no other constraining information is known so that local connectivity i.e., node strength as measured by degree, are primarily important. In the case of spatially embedded networks, however, this choice generally yields spurious communities due to the lack of mixing and the strong spatial clustering of the network Gilarranz2020SciRep . A number of works have introduced and studied alternative null models designed specifically for spatially embedded network, for example the gravity model Expert2011PNAS , the radiation model Simini2012Nature ; Sarzynska2016CompNet , and the exponential decay model Cerina2012PLOSOne , all of which similarly combine local nodal properties with information regarding the explicit distance between nodes. The gravity model, for instance Expert2011PNAS , uses a null model given by $P_{ij}=k_{i}k_{j}f(d_{ij})/Z$, where $d_{ij}=d(x_{i},x_{j})$ is the Euclidean distance between nodes $i$ and $j$, $f(d)$ is function that decreases monotonically as the distance $d$ increases, and $Z$ is a normalization constant set to ensure that $\sum_{i=1}^{N}\sum_{j=1}^{N}P_{ij}=\sum_{i=1}^{N}\sum_{j=1}^{N}A_{ij}$. In what follows we will use the gravity model with the choice $f(d_{ij})=\text{exp}(-d_{ij}^{2})$, i.e., a Gaussian curve. Thus, these null models predict expected values of links in a way that balances the nodes’ overall connectivity with their proximity in space.
@@
Using modularity to partition large ensembles of individual units into multiple coherent groups has been used in a wide range of applications Red2011SIAMRev ; Porter2005PNAS ; Porter2007PhysA ; Zhang2008PhysA ; Macon2012PhysA ; Bassett2011PNAS ; Wymbs2012Neuron ; Yin2020PNAS ; Parker2015ACS ; Billock2020Sex . A wide range of other partitioning techniques exist as well, for instance, Laplacian eigenmaps Belkin2003Neur , diffusion maps Coifman2005PNAS , $k$-means Lloyd1982 , Otsu’s method Otsu1979 , but in contrast to these methods the modularity method has the added component of assigning to any partition a measure of the quality of a given partition, i.e., the actual modularity value $Q$. However, for the task of identifying subpopulations that correspond to multimodality in data, existing methods fail. Specifically, community detection using existing null models, as well as the other methods referred to above, make strong partitions that do not correspond to bimodality. Take, for instance, a data set $\{x_{i}\}_{i=1}^{N}$ where each data point $x_{i}$ is identically and independently drawn from the following superposition of Normal distributions:
@@
Note that for the case of evenly-split clusters, $\eta=1/2$, equal variances, $\sigma_{1}^{2}=\sigma_{2}^{2}=\sigma^{2}$, and opposite means $\mu_{1}=-\mu_{2}=\mu$, this underlying distribution is bimodal if and only if $\mu>\sigma$. Thus, for the case of $\eta=1/2$, $\mu=1/2$, and $\sigma=1$ the underlying distribution is clearly unimodal. However, as an example of the failure of existing methods for this task and motivation for a new null model, we consider a data set of size $N=1000$ drawn from this choice of distribution and examine the optimal community partitions using both the classical NG null model Newman2004PRE as well as the gravity null model Cerina2012PLOSOne , where the underlying network is generated using $r=1/2$. In Fig. 1 we plot in panel (a) the distribution of the drawn data, with the two clusters identified using the NG null model in blue circles and red triangles (with the underlying distribution plotted in dashed black) and in panel (b) a heat map of the corresponding modularity matrix with the cut location indicated by black lines. In panels (c) and (d) we plot similar results for the gravity model, using $P_{ij}\propto k_{i}k_{j}\text{exp}(-|x_{i}-x_{j}|)$. Note that both methods make a strong partition near $x\approx 0$ (the modularity for the NG and gravity cases were optimized at $Q=0.3705$ and $0.1814$, respectively), despite the clear lack of any bimodality in the data. These results are similar to those obtained by using Laplacian eigenmaps, diffusion maps, $k$-means (with $k=2$, and Otsu’s method, all of which tend to favor a partition in the data near the middle, regardless of whether there is any bimodality in the underlying distribution. This failure of existing partitioning methods to faithfully correspond to multimodality thus motivates the need for a new null model that addresses this problem.
@@
III A new proximity-based null model
@@
IV Synthetic data benchmarks
@@
Moving to the modularity heat map for $\mu=1.2$ in Fig. 3(g), we now see the emergence of a robust range of $r$ where a local maxima in $Q$ exists, near $x_{\text{cut}}=0$ between roughly $r=0.4$ and $1.0$. Choosing $r=0.8$ we see in panel (h) the bands of the modularity matrix now organize into coherent sections of positive and negative values, resulting in a stronger partition of the data, as illustrated in the two clusters in panel (i). The partition of data and identification of two subpopulations becomes stronger as $\mu$ is increased, as illustrated in Fig. 4, with more robust bands of $r$ for which the heat maps of $Q$ yields a significant local maximum [panels (a), (d) and (g)] stronger bands of positive and negative vales in the modularity matrix [panels (b), (e), and (h)], and strong, accurate partitions of the data [panels (c), (f), and (i)]. Overall, using the new spatially informed null model successfully identifies subpopulations when the underlying data is bimodal ($\mu=1.2$, $1.5$, $1.8$, and $2.0$) and accurately does not when the underlying data is unimodal ($\mu=0.8$ and $1.0$). In addition to agreeing with the unimodality or bimodality of the known underlying distribution for these synthetic benchmark examples, we also report for each case the likelihood $p_{\text{dip}}$ of rejecting a null hypothesis of the data being unimodal using Hartigan’s dip test Hartigan1985 : for $\mu=0.8$, $1.0$, $1.2$, $1.5$, $1.8$, and $2.0$ we find, respectively $p_{\text{dip}}=0.9886$, $0.6166$, $3.6\times 10^{-3}$, $<1.0\times 10^{-3}$, $<1.0\times 10^{-3}$, and $<1.0\times 10^{-3}$, which match up well with our results.
@@
V Identifying cell states using oxidative stress

### v3

@@ Abstract
Detecting communities in large complex networks has found a wide range of applications in physical, biological, and social sciences by identifying mesoscopic groups based on the links between individual units. Moreover, community detection approaches have been generalized to various data analysis tasks by constructing networks whose links depend on individual units’ measurements. However, identifying well-separated subpopulations in data sets, e.g., multimodality, still presents challenges as both community detection with existing null models and other partition methods either fail to give partitions that correspond to dips in the data or give partitions that do not correspond to dips in the data. Here we introduce a new spatially informed null model for this task that takes into account spatial structure but does not explicitly depend on distances between measurements. We find that community detection using this null model successfully identifies subpopulations in multimodal data and accurately does not for unimodal data. This new method represents a complement to statistical methods, as we treat data directly using a network science approach. We apply this new null model to first distinguish interphase and mitotic cell cycle phases and then S and G2 cell cycle phases in a group of Dictyostelium discoideum cells using measurements of oxidative stress, which have been shown to correlate strongly with cell-cycle behaviors.
keywords:
Bimodality, Community Detection, Cell Cycle, Oxidative Stress
1 Introduction
@@
Community detection in complex networks has found applications in a wide range of applications by identifying functional and mesoscopic groups based on patterns of interactions and relationships between individual units [1, 2, 3, 4]. Examples include groupings in online social networks [5], partnerships in political institutions [6, 7, 8, 9], architectures of neuronal organization [10, 11, 12], identifying antibiotic resistance in patients [13], and analysis of contact networks for sexually-transmitted diseases [14]. In principle, these methods for identifying groups can be generalized to quantitative data sets by constructing networks from the data where nodes correspond to measurements and links are placed based on the pair-wise distances between measurements. These generalized methods for network-based data analysis, however, have their drawbacks, as community structures generically emerge in spatially embedded networks even when the data shows no signs of having significant subpopulations, i.e., multimodality or simply making erroneous partitions that do not correspond with any significant divide in the data [15]. Thus, developing community detection frameworks that correspond more precisely with the identification of subpopulations, i.e., multimodality, and likewise accurately failing to do so in the absence of subpopulations, is an important task, as testing for multimodality in real datasets remains a challenging problem with few robust methodologies [16].
@@
One such problems is the classification of cell cycle phases using measurements of oxidative stress. Reactive oxygen species are hypothesized to play important roles in microbial development of cells [17, 18]. Reactive oxygen species are also linked with cell cycles, with mitotic cells having been shown to exhibit higher levels of oxidative stress than cells in interphase. Thus, subpopulations of cells distinguished by oxidative stress may correspond to different cell cycles. Given the practical challenges of taking individual cell measurements, linking cell cycle phases with measurements of oxidative stress could be useful in synthetic engineering tasks such as synchronizing cell cycles as well as developing a better understanding of the sources of heterogeneity between individual cells and throughout entire groups [19]. Here we seek to identify significant subpopulations in a group of Dictyostelium discoideum cells using oxidative stress measurements quantified by the log-ratio of dihydrodichlorofluorescein diacetate (DCF) to carboxyfluorescein diacetate (CF). Technical details for the experimental measurments and data collection are provided in the Appendix.
@@
To identify subpopulations in this and other datasets we utilize the modularity method for detecting communities in networks [20]. This method is based on comparing an actual network structure to an appropriate null model that quantifies the expectation of the existence and strength of each pairwise link. Rather than using the classical Newman-Girvan null model [20] or other null models that incorporates spatial structure by explicitly using distances [21, 22, 23, 24], we introduce a new null model that is spatially informed but does not explicitly use distances. Thus, similar to the Newman-Girvan null model, this new null model uses primarily the local connectivity properties of nodes, but like distance-dependent null models a strong spatial structure is assumed without strong mixing. In essence, the null model is designed to assume that a node is most likely connected to the number of nodes matching its degree that are all close by, but this likelihood is not explicitly a function of any distance. We test this new null model on a series of synthetic data benchmarks to examine the effects of separation and asymmetry in the data, finding that it successfully identifies subpopulations, i.e., multimodality, when they are present and accurately fails to do so when they are not, i.e., the data is unimodal. We then turn to the task of partitioning cells according to their cell cycle phases using measurements of oxidative stress. Throughout our results, we will present the results from Hartigan’s dip test [16] as a point of comparison, but emphasize here that our new method should be viewed as a complement to this statistical test, as we aim to treat and make hypotheses about data directly.
@@
The remainder of this paper is organized as follows. In Sec. 2 we describe network-based approaches to identifying subpopulations in data using modularity and discuss the choice of null models. We then introduce a new spatially informed null model designed for identifying multimodality in data. In Sec. 4 we evaluate this method with our new null model over a collection of synthetic data benchmarks. In Sec. 5 we apply our method to real data taken from an ensemble of Dictyostelium discoideum cells. Specifically, we use oxidative stress measurements first to identify cells in interphase vs mitosis, then in S and G2 phases. In Sec. 6 we conclude with a discussion of our results.
@@
2 Data analysis, community detection, and the failure of existing null models
@@
Motivated by the task of identifying cell-cycle phases in experimental data of oxidative stress in Dictyostelium discoideum cells, we consider data sets $\{x_{i}\}_{i=1}^{N}$ consisting of $N$ measurements with $i=1,\dots,N$. In our motivating example, each $x_{i}$ represents a relative measurement of oxidative stress in a cell, which is strongly correlated with the phase (interphase or mitosis) of the cell cycle. (More details on the data are given in Sec. 5 and in Appendix A) In general we assume that each data point is a scalar, i.e., $x_{i}\in\mathbb{R}$, and therefore can be sorted in ascending order so that $x_{1}\leq x_{2}\leq\cdots\leq x_{N}$. We then construct from this data set a network of $N$ nodes where each node $i$ corresponds to a measurement $x_{i}$ and links are placed between nodes based on the distance between corresponding measurements. Many choices can be reasonably made in terms of how links are placed, but here we consider the simple choice where a link is placed between nodes $i$ and $j$ if the measurements $x_{i}$ and $x_{j}$ satisfy $|x_{i}-x_{j}|\leq rs_{x}$, where $r$ is a tunable local connectivity parameter and $s_{x}$ is the standard deviation of the dataset $\{x_{i}\}_{i=1}^{N}$. This network is described by the adjacency matrix $A$, whose entries $A_{ij}=1$ if a link exists between nodes $i$ and $j$, and otherwise $A_{ij}=0$.
@@
We then seek to map the task of identifying subpopulations in the data, i.e., multimodality, to finding communities in the corresponding network. Community detection has long been a topic of interest in the complex networks community and has a rich literature [1, 2, 4]. Here we choose the modularity method [3], where a given network structure is compared to an appropriately chosen null model and then partitioned into clusters that are more highly connected than expected by maximizing the so-called modularity. Mathematically, the modularity $Q$ can be expressed as
@@
Instead, our challenge here is in finding a null model (encapsulated in the matrix $P$) that allows us to find communities that accurately map to subpopulations, i.e., multimodality, in the original data. The classical choice for a null model was first presented by Newman and Girvan [20] and is given by the expectation of a link being proportional to the product of the node-wise degrees, i.e., $P_{ij}=k_{i}k_{j}/(2M)$, where $k_{i}=\sum_{j=1}^{N}A_{ij}$ is the degree of node $i$. This null model is very effective when the network structure is well mixed and no other constraining information is known so that local connectivity i.e., node strength as measured by degree, are primarily important. In the case of spatially embedded networks, however, this choice generally yields spurious communities due to the lack of mixing and the strong spatial clustering of the network [15]. A number of works have introduced and studied alternative null models designed specifically for spatially embedded network, for example the gravity model [21], the radiation model [22, 23], and the exponential decay model [24], all of which similarly combine local nodal properties with information regarding the explicit distance between nodes. The gravity model, for instance [21], uses a null model given by $P_{ij}=k_{i}k_{j}f(d_{ij})/Z$, where $d_{ij}=d(x_{i},x_{j})$ is the Euclidean distance between nodes $i$ and $j$, $f(d)$ is function that decreases monotonically as the distance $d$ increases, and $Z$ is a normalization constant set to ensure that $\sum_{i=1}^{N}\sum_{j=1}^{N}P_{ij}=\sum_{i=1}^{N}\sum_{j=1}^{N}A_{ij}$. In what follows we will use the gravity model with the choice $f(d_{ij})=\text{exp}(-d_{ij}^{2})$, i.e., a Gaussian curve. Thus, these null models predict expected values of links in a way that balances the nodes’ overall connectivity with their proximity in space.
@@
Using modularity to partition large ensembles of individual units into multiple coherent groups has been used in a wide range of applications [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. A wide range of other partitioning techniques exist as well, for instance, Laplacian eigenmaps [25], diffusion maps [26], $k$-means [27], Otsu’s method [28], but in contrast to these methods the modularity method has the added component of assigning to any partition a measure of the quality of a given partition, i.e., the actual modularity value $Q$. However, for the task of identifying subpopulations that correspond to multimodality in data, existing methods fail. Specifically, community detection using existing null models, as well as the other methods referred to above, make strong partitions that do not correspond to bimodality. Take, for instance, a data set $\{x_{i}\}_{i=1}^{N}$ where each data point $x_{i}$ is identically and independently drawn from the following superposition of Normal distributions:
@@
Note that for the case of evenly-split clusters, $\eta=1/2$, equal variances, $\sigma_{1}^{2}=\sigma_{2}^{2}=\sigma^{2}$, and opposite means $\mu_{1}=-\mu_{2}=\mu$, this underlying distribution is bimodal if and only if $\mu>\sigma$. Thus, for the case of $\eta=1/2$, $\mu=1/2$, and $\sigma=1$ the underlying distribution is clearly unimodal. However, as an example of the failure of existing methods for this task and motivation for a new null model, we consider a data set of size $N=1000$ drawn from this choice of distribution and examine the optimal community partitions using both the classical NG null model [20] as well as the gravity null model [24], where the underlying network is generated using $r=1/2$. In Fig. 1 we plot in panel (a) the distribution of the drawn data, with the two clusters identified using the NG null model in blue circles and red triangles (with the underlying distribution plotted in dashed black) and in panel (b) a heat map of the corresponding modularity matrix with the cut location indicated by black lines. In panels (c) and (d) we plot similar results for the gravity model, using $P_{ij}\propto k_{i}k_{j}\text{exp}(-|x_{i}-x_{j}|)$. Note that both methods make a strong partition near $x\approx 0$ (the modularity for the NG and gravity cases were optimized at $Q=0.3705$ and $0.1814$, respectively), despite the clear lack of any bimodality in the data. These results are similar to those obtained by using Laplacian eigenmaps, diffusion maps, $k$-means (with $k=2$, and Otsu’s method, all of which tend to favor a partition in the data near the middle, regardless of whether there is any bimodality in the underlying distribution. This failure of existing partitioning methods to faithfully correspond to multimodality thus motivates the need for a new null model that addresses this problem.
@@
3 A new proximity-based null model
@@
4 Synthetic data benchmarks
@@
Moving to the modularity heat map for $\mu=1.2$ in Fig. 3(g), we now see the emergence of a robust range of $r$ where a local maxima in $Q$ exists, near $x_{\text{cut}}=0$ between roughly $r=0.4$ and $1.0$. Choosing $r=0.8$ we see in panel (h) the bands of the modularity matrix now organize into coherent sections of positive and negative values, resulting in a stronger partition of the data, as illustrated in the two clusters in panel (i). The partition of data and identification of two subpopulations becomes stronger as $\mu$ is increased, as illustrated in Fig. 4, with more robust bands of $r$ for which the heat maps of $Q$ yields a significant local maximum [panels (a), (d) and (g)] stronger bands of positive and negative vales in the modularity matrix [panels (b), (e), and (h)], and strong, accurate partitions of the data [panels (c), (f), and (i)]. Overall, using the new spatially informed null model successfully identifies subpopulations when the underlying data is bimodal ($\mu=1.2$, $1.5$, $1.8$, and $2.0$) and accurately does not when the underlying data is unimodal ($\mu=0.8$ and $1.0$). In addition to agreeing with the unimodality or bimodality of the known underlying distribution for these synthetic benchmark examples, we also report for each case the likelihood $p_{\text{dip}}$ of rejecting a null hypothesis of the data being unimodal using Hartigan’s dip test [16]: for $\mu=0.8$, $1.0$, $1.2$, $1.5$, $1.8$, and $2.0$ we find, respectively $p_{\text{dip}}=0.9886$, $0.6166$, $3.6\times 10^{-3}$, $<1.0\times 10^{-3}$, $<1.0\times 10^{-3}$, and $<1.0\times 10^{-3}$, which match up well with our results.
@@
5 Identifying cell states using oxidative stress